# persiansort : an alternative to mergesort inspired by persian rug

### v1

@@ 3 New Algorithm Design
In persiansort, warping (divide step) can transition from wp=4 to wp=n+1 where ”n” is the size of the original dataset. Alternative functions for warping may also be utilized, for instance: .
@@
The most crucial aspect is knotting many components of this algorithm, which can use of runs. To this end, we designate a pointer named i in line 18. Consequently, in each division, we retain the initial memory as pre-warp and by employing two knots in lines 19 and 22 the size of this segment (pre-warp) expands based on the data layout. The division process transpires subsequent to the operation of this knot, which can detect previously sorted data in various segments of the division. If the length of the sorted data exceeds wp, the second knot utilizes all of the runs. It also finds a substantial portion of the ordered data with a length inferior to wp.
@@
Additional knots may be incorporated into the persiansort algorithm as required, prior to line number 16, enabling hybridization with insertion sort when the data count falls below a specified threshold. Preliminary testing findings indicate that the application of this knot enhances speed in random data by around 10 percent. In addition to these, we can change of the max-jump during warping.
Line 24 governs the final stage of copying to the auxiliary array.
Line 25 determines whether we are transferring data from the main array to the auxiliary array or from the auxiliary array to itself.
Lines 39 and 40 quantify the auxiliary memory and allocate it accordingly.

### v2

@@ 3 New Algorithm Design
In persiansort, warping (divide step) can transition from wp=4 to wp=n+1 where ”n” is the size of the original dataset. Alternative functions for warping may also be utilized, for instance: max-jump= .
@@
The most crucial aspect is knotting many components of this algorithm, which can use of runs. To this end, we designate a pointer named i in line 21. Consequently, in each division, we retain the initial memory as pre-warp and by employing two knots in lines 22 and 25 the size of this segment (pre-warp) expands based on the data layout. The division process transpires subsequent to the operation of this knot, which can detect previously sorted data in various segments of the division. If the length of the sorted data exceeds wp, the second knot utilizes all of the runs. It also finds a substantial portion of the ordered data with a length inferior to wp.
@@
Additional knots may be incorporated into the persiansort algorithm as required, prior to line number 20, enabling hybridization with insertion sort when the data count falls below a specified threshold. Preliminary testing findings indicate that the application of this knot enhances speed in random data by around 10 percent. In addition to these, we can change of the max-jump during warping.
Line 27 governs the final stage of copying to the auxiliary array.
Line 28 determines whether we are transferring data from the main array to the auxiliary array or from the auxiliary array to itself.
Lines 42 and 43 quantify the auxiliary memory and allocate it accordingly.